OpenStudy (anonymous):

aprrox tan(-.1) using Taylor polynomials with n=3

OpenStudy (anonymous):

whew only 3 is good \[\tan(x)=x+\frac{x^3}{3}+\frac{2x^5}{3\times 5}+...\]

OpenStudy (anonymous):

put \(x=-.1\) and a calculator, which is really kind of dumb since you can just put \(\tan(-.1)\) directly in to a calculator

OpenStudy (anonymous):

Thanks! Can I ask ya another quick question? My next problem says eval sqrt 101 using taylor. Do I just call it sqrt 100 and x=1?

OpenStudy (anonymous):

use taylor polynomail witj n=3 to approx. (101)^1/2

OpenStudy (anonymous):

yes i think so

OpenStudy (anonymous):

use the taylor polynomial for \[\sqrt{100+x}\]

OpenStudy (anonymous):

or x=101 a=0?

OpenStudy (anonymous):

use \[f(x)=\sqrt{100+x}\] because you want to expand around \(x=0\) then you can put \(x=1\)

OpenStudy (anonymous):

how many terms you need?

OpenStudy (anonymous):

I think my books answer has a=100? 3 terms

OpenStudy (anonymous):

the firt term of answer is 10 hen 1/20 (x-100) .......

OpenStudy (anonymous):

ok good expand about zero get \[10+\frac{x}{20}\]

OpenStudy (anonymous):

should be \(x\) up top, not \(100-x\)

OpenStudy (anonymous):

im with ya

OpenStudy (anonymous):

you are expanding about 0, the 100 is in the function

OpenStudy (anonymous):

i think the next term is \(-\frac{x^2}{800}\) but it is late so i could be wrong

OpenStudy (anonymous):

oh yeah wrong! i am off by a zero

OpenStudy (anonymous):

thats correct......I managed to get a little lost with it though ); Can I check the derivatives with you?

OpenStudy (anonymous):

I can start a new post if youd like for another medal?

OpenStudy (anonymous):

sure but it comes with no guarantees

OpenStudy (anonymous):

lol no i don't need any more thank you i got this bike for reaching 1000, after that i can't seem to redeem them anywhere

OpenStudy (anonymous):

lol cool

OpenStudy (anonymous):

ok so from the biginning ( sorry) werr going to say f(x)= (100+x)^1/2 right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

also a=0

OpenStudy (anonymous):

yes, aka mclauren

OpenStudy (anonymous):

so they say to approx with n=3 so.......ned first 3 derivs.

OpenStudy (anonymous):

f'(x)= 1/2(100+x)^1/2

OpenStudy (anonymous):

exponent is \(-\frac{1}{2}\) but it is easier to write \[f'(x)=\frac{1}{2\sqrt{100+x}}\] you have to evaluate at \(x=0\) so want a the radical form probably

OpenStudy (anonymous):

although the form \[\frac{1}{2}(100+x)^{-\frac{1}{2}}\] is good for taking the second derivative

OpenStudy (anonymous):

\[f \prime \prime (x) + -1/4\sqrt{(100+x)^3}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

and so \(f''(0)=-\frac{1}{4\times 10^3}=-\frac{1}{4000}\)

OpenStudy (anonymous):

and \(\frac{f''(0)}{2}=-\frac{1}{8000}\) so you are good to go with three terms

OpenStudy (anonymous):

why divide by 2?

OpenStudy (anonymous):

ahhhh

OpenStudy (anonymous):

2!